OpenStudy (anonymous):

Given f(x) = 2x -1 and g(x) = 1- x^2 Find g(g(x))

OpenStudy (anonymous):

Do you mean f(g(x))?

OpenStudy (anonymous):

no it says it like that

OpenStudy (anonymous):

g(g(x))=1-g(x)^2=1-(1-x^2)^2